OpenStudy (anonymous):

Luke mowed a lawn today to earn some extra money. He spent $2.10 on fuel and $4.45 on lunch. He charged $13.50 to mow the lawn. What was his profit for mowing the lawn? A. $20.05 B. $15.85 C. $11.15 D. $6.95

OpenStudy (anonymous):

help i will fan and medal @imtotallystuck @17Trinity17

OpenStudy (anonymous):

$6.95 in profit

OpenStudy (idealist10):

13.5-2.1-4.45=11.4-4.45=6.95 Answer: 6.95

OpenStudy (anonymous):

Ok, Thank You Guys So Much!

OpenStudy (idealist10):

You're welcome. The answer is D.

OpenStudy (anonymous):

.The table shows the low outside temperatures for Monday, Tuesday, and Wednesday. By how many degrees did the low temperature increase from Monday to Tuesday?

OpenStudy (anonymous):

nvm